OpenStudy (anonymous):

Solve x 5 * 18^(6x)=26

OpenStudy (anonymous):

\[\huge 90^{6x}=26\] Is that the correct first step?

OpenStudy (isaiah.feynman):

Nope. Can't multiply 5 to 18, they have different exponents. Instead divide both side by 5 so 18 is alone.

OpenStudy (anonymous):

thats the naswer

OpenStudy (anonymous):

@abswag13 I can type things into mathway too. I'm here to learn. Not looking for straight answers @Isaiah.Feynman so I have \[\huge 18^{6x}=\frac{ 26 }{ 5 }\]

OpenStudy (anonymous):

i WAS going to explain if you waited a minute more but you didn't

OpenStudy (anonymous):

snappy little flutter

OpenStudy (isaiah.feynman):

Okay. Rewrite 18 as 3 times 6 so its easier.

OpenStudy (anonymous):

@abswag13 I wasn't trying to sound snappy or rude. Sorry if I came off that way. I would prefer the work explained before given any answer.

OpenStudy (anonymous):

what ever >:(

OpenStudy (anonymous):

So I have\[\huge (3*6)^{6x} = \frac{ 26 }{ 5 }\]

OpenStudy (isaiah.feynman):

Never mind that. Take the natural logarithm of both sides.

OpenStudy (isaiah.feynman):